OpenStudy (ars22):

How do you reduce the following to lowest rooms and then to mixed expressions? x^2 - 16 -------- x (x - 4)

sam (.sam.):

\[\frac{(x^2 - 16)}{x(x-4)}\] Factor \[ =\ \frac{(x-4)(x+4)}{x(x-4)}\] \[=\frac{\cancel{(x-4)}(x+4)}{x\cancel{(x-4)}}\] =?